OpenStudy (anonymous):

The derivative of f(x)=-(x-1)^2-2 is... f'(x)= -x^2 + 2x - 3 Now find the slope.

OpenStudy (anonymous):

The derivative is the slope...so the slope is f'(x)= -x^2+2x-3 If you wanna find slope at certain point then replace that point in f'(x) for example if you wanna find slope at x=1 then f'(1) = -(1)^2 +2(1) -3 = -1+2-3 = -4 + 2 = -2 so slope at point x=1 is -2...you can find slope at any given point by just replacing x by that number in f'(x) or at its derivative. Remember derivative is the slope

OpenStudy (anonymous):

got it?